# Adaptive Incentive Design with Learning Agents

### v1

@@ 1 Introduction
Incentive mechanisms play a crucial role in many societal systems, where the outcomes are governed by the selfish interactions of a large number users (or algorithms on their behalf) and the system operator. The outcome of such strategic interactions, characterized by the Nash equilibrium, is often sub-optimal because of the fact that individual agents often do not account for the externality of their actions (i.e., how their actions affect the costs of others), when minimizing their own costs. An important way to address the sub-optimality is to provide agents with incentives that align their individual goal of cost minimization with the goal of minimizing the total cost of the societal system[Lev85, HLO82, Baş84].
The system operator faces two key challenges when designing such incentive mechanisms. First, due to the system’s scale and privacy concerns, the designer often has limited information about individual agents’ costs and the cost function of the entire system. Second, the system operator often needs to consider the learning behavior of agents, who repeatedly update their strategies in response to the incentive mechanism, especially when the physical system experiences a random shock and players are trying to reach a new equilibrium [BG14, CM21, MKWS21]. As a result, the designer cannot infer the equilibrium to predict agents’ behavior, making incentive design a challenging task. This leads to the following question:
@@
We provide sufficient conditions on the underlying game such that the set of fixed points is a singleton, making the socially optimal incentive mechanism unique (Proposition 3.1). We also characterize sufficient conditions that guarantee the convergence of the coupled dynamical system (Proposition 3.3). Since the convergent strategy profile and incentive mechanism correspond to a fixed point that is also socially optimal, these sufficient conditions ensure that the dynamic incentive mechanism eventually induces a socially optimal outcome in the long run. To prove this result, we leverage the timescale separation between the strategy update and the incentive update, using results from the theory of two-timescale dynamical systems [Bor97]. This theory allows us to analyze the convergence of the coupled dynamical system in two steps by decoupling the convergence of the strategy update and the incentive mechanism update. First, the convergence of the strategy update (on the faster timescale) is analyzed by viewing the incentive mechanism (on the slower timescale) as static. In our work, we off-load this convergence to the rich literature on learning in games (e.g., [FL98, San10, MS96]). Second, the convergence of the incentive mechanism update is analyzed via the associated continuous-time dynamical system, evaluated at the converged value of the strategy update (i.e., Nash equilibrium). The sufficient conditions developed in Proposition 3.3 are built on results on the convergence of non-linear dynamical systems from cooperative dynamical systems theory [Hir85] and Lyapunov-based methods [Sas13].
@@ 1.1 Related Works
Two-timescale Learning Dynamics: Learning dynamics where incentives are updated at a slower timescale than the strategy of agents have been studied in [MNPQ22, COMP23, PBMT17, OP22, RF20, LLY+21, LNL+23]. Specifically, [MNPQ22] studies Stackelberg games with a single leader and a population of followers. In their approach the leader deploys gradient-based update and followers update their strategies using replicator dynamics. [COMP23] and [PBMT17] consider incentive design in affine congestion games, where the incentives are updated using a distributed version of gradient descent. [OP22] studies incentive design to control traffic congestion on a single highway using gradient based incentive updates. [RF20] studies the problem of incentive design while learning the cost functions of agents. The authors assume that both the cost functions and incentive policies are linearly parameterized, and the incentive updates rely on the knowledge of agents’ strategy update rules instead of just the current strategy as in our setting. Additionally, [LLY+21, LNL+23] propose a two-timescale discrete-time learning dynamics, where agents adopt a mirror descent-based strategy update, and the social planner updates the incentive parameter according to a gradient based method. The convergence of these gradient-based learning dynamics relies on the assumption that the social cost given agents’ equilibrium strategy is convex in the incentive parameter. However, the convexity assumption is restrictive as highlighted by an example in Appendix A.
@@
Same-timescale Learning Dynamics: The problem of steering non-cooperative agents to a desired Nash equilibrium by considering an incentive update which runs on the same timescale as strategy updates is studied in [SCM22a, SCM22b, ZFA+23]. Particularly, in [SCM22a] the authors study such updates in a specific setting of quadratic aggregative games. In [SCM22b], the authors consider a setup where the agents’ cost is only dependent on their own action and the price signal which is provided by operator which is used by agents based on their individual reliability over the operator. The aim is to design this price signal so that the reliability on the price signal is retained while steering the equilibrium to the desired point. In [ZFA+23], the authors consider the problem of steering no-regret learning agents to optimal equilibrium but the computation of incentive mechanism requires solving an optimization problem at each time step.
@@
Learning in Stackelberg Games: Our work is also related to the literature on learning in Stackelberg games, where the planner often has limited information about the game between agents and need to design optimal mechanism by dynamically receiving feedback from agents’ responses (for instance, refer to [BHP14, LCM09, PSTZ19, BJWX21, CFD24, JGA+22, MSRM23]). Typically, this line of research posit structural assumptions on the game between followers (e.g., finite action space of agents, or linearly parametrized utility function of agents)[BHP14, LCM09, PSTZ19, BJWX21, CFD24, JGA+22], or only ensures convergence to a locally optimal solution [MSRM23].
@@ 2.1.1 Atomic Games
Under Assumption 2.1, it is known that, for any $p\in\mathbb{R}^{|\mathcal{I}|}$, the Nash equilibrium $x^{\ast}(p)$ exists, is unique and is Lipscthiz continuous in $p$ [Daf88]. Moreover, the socially optimal strategy $x^{\dagger}$ is also unique.
@@ 2.1.2 Non-atomic Games
Under Assumption 2.2, Nash equilibrium $x^{\ast}(p)$ exists, is unique and Lipscthiz continuous for every $p\in\mathbb{R}^{|\tilde{\mathcal{I}}|}$ [San10, Daf88]. Moreover, the socially optimal strategy distribution is also unique.
@@ 2.2 Coupled Strategy and Incentive Update
    Equilibrium update ([CFD24, JGA+22]): The strategy update incorporates a Nash equilibrium strategy profile with respect to the incentive mechanism in step $k$:
@@
    Best response update ([FL98, HS02]): The strategy update incorporates a best response strategy with respect to the strategy and the incentive mechanism in step $k$:
@@
    Gradient-based update ([LW94, LM13, San10]): The strategy update incorporates a smoothed strategy update with respect to the strategy and the incentive mechanism in step $k$:
@@
    where $z_{i}(x_{k},p_{k})=x_{k}-\eta\nabla_{x_{i}}c_{i}(x_{k},p_{k})$, $\eta$ is step size, and $h(\cdot),\tilde{h}(\cdot)$ are regularizers. If $h(\cdot)$ is a quadratic function then the update becomes projected gradient descent [MZ19]. Furthermore, if $\tilde{h}(\cdot)$ is the entropy function then the update becomes a perturbed best-response update[San10].
@@
Our externality-based incentive update contrasts with the gradient-based incentive update considered in the literature [LLY+21]. Such a gradient-based incentive update typically requires that the equilibrium social cost function $\Phi(x^{\ast}(p))$ (resp. $\tilde{\Phi}(\tilde{x}^{\ast}(\tilde{p}))$) be strongly convex in the incentive mechanism $p$ (resp. $\tilde{p}$). However, this requirement is too restrictive and is not satisfied even in simple games due to the fact that the equilibrium strategy vector is typically non-convex in the incentive mechanism. In Appendix A, we provide an example of a two-link routing game, where the equilibrium social cost function is non-convex in the incentive mechanism, and the gradient-based learning dynamics do not converge.
@@ 3.1 Fixed point analysis
First, we show that $P^{\dagger}$ is non-empty. That is, there exists $p^{\dagger}$ such that $e(x^{\ast}(p^{\dagger}))=p^{\dagger}$. Define a function $\theta(p)=e(x^{\ast}(p))$. We note that Assumption 2.1 guarantees that $\theta$ is well defined. Thus the problem is reduced to showing existence of fixed point of $\theta(\cdot)$, for which we use the Schauder fixed point theorem [Sma80].
@@
where $J(x,p)$ is the game Jacobian defined as $J_{i}(x,p)=\nabla_{{x}_{i}}\ell_{i}(x)+p_{i}$ for every $i\in\mathcal{I}$. Next, under Assumption 2.1, the variational inequality characterization of Nash equilibrium [FP07], we know that $x^{\ast}(p^{\dagger})$ is a Nash equilibrium if and only if
@@
Under Assumption 2.2, the variational inequality characterization of Nash equilibrium for non-atomic games [San10], we know that ${\tilde{x}^{\ast}}(\tilde{p}^{\dagger})$ is a Nash equilibrium if and only if
@@ 3.2 Convergence to optimal incentive mechanism
In Assumption 3.1, (i) is a standard assumption on step-sizes that allows us to analyze the convergence of the discrete-time learning updates via a continuous-time differential equation [Bor97]. Additionally, (ii) assumes that the incentive update occurs at a slower timescale compared to the update of strategies.
In view of Assumption 3.1-(ii), it is necessary that the strategy updates converge if the incentives are held fixed. This condition holds true for a variety of updates commonly studied in learning in games literature. For instance, the equilibrium update (cf. (4)) converges to Nash equilibrium in one-step. Similarly, the best response update (cf. (5)) converges in zero sum games [Har98], potential games [SMK18], and dominance solvable games [Nac90]. The gradient-based update (cf. (6)) converge in concave games [Ros65]. Therefore, in this work we do not focus on studying the convergence of strategy updates with fixed incentive mechanisms and we pick any off-the-shelf convergent strategy update. Our goal is to characterize conditions under which the coupled strategy and incentive updates ($x$-update)-($p$-update) and ($\tilde{x}$-update)-($\tilde{p}$-update) converge. Therefore, we impose the following assumption which is sufficient to ensure convergence of strategy updates ($x$-update) and ($\tilde{x}$-update) with fixed incentive mechanisms. Particularly, we assume that if the incentive mechanism is fixed, then the corresponding continuous-time dynamical system converges to the Nash equilibrium associated with that incentive mechanism.
@@
Using the standard approximation arguments from [Bor09] it is evident that Assumption 3.1 and 3.2 are sufficient to ensure convergence of strategy updates ($x$-update) and ($\tilde{x}$-update) with fixed incentive mechanisms.
@@
satisfy [^2] The results in this paper hold even if the updates ($x$-update)-($p$-update) and ($\tilde{x}$-update)-($\tilde{p}$-update) are perturbed with martingale difference noise [BP18].
@@
Under Assumption 3.1, the two-timescale approximation theory [Bor97, BP18] suggests that the strategy update ($x$-update) is a fast transient while the incentive update ($p$-update) is a slow component. This separation in timescales allow us to study the convergence of these updates in two stages. First, we study the convergence of fast strategy updates, for every fixed value of incentive. Second, we study the convergence of slow incentive updates, assuming that the fast strategy updates have converged to the equilibrium.
@@
Since $\sup_{k\in\mathbb{N}}(\|x_{k}\|+\|p_{k}\|)<+\infty$ and $\lim_{k\rightarrow\infty}\beta_{k}/\gamma_{k}=0$ (cf. Assumption 3.1), the term $\frac{\beta_{k}}{\gamma_{k}}\left(e(x_{k})-p_{k}\right)$ in (20) goes to zero as $k\rightarrow\infty$. Consequently, leveraging the standard approximation arguments [Bor09, Lemma 1, Section 2.2], we conclude that the asymptotic behavior of the updates in (20) is same as that of the following dynamical system
@@
Note that $p^{\dagger}$ is the fixed point of the trajectories of dynamical system (23) (cf. Proposition 3.1). Conditions (C1) and (C2) in Proposition 3.3 are based on results from non-linear dynamical systems which ensure convergence of (23). In particular, (C1)-(i) (resp. (C1)-(ii)) builds on the cooperative dynamical systems theory [Hir85], which ensures that $\mathbb{R}_{+}^{|\mathcal{I}|}$ (resp. $\mathbb{R}_{-}^{|\mathcal{I}|}$) is positively invariant for (23) and $p^{\dagger}\in\mathbb{R}_{+}^{|\mathcal{I}|}$ (resp. $p^{\dagger}\in\mathbb{R}_{-}^{|\mathcal{I}|}$) is asymptotically stable. On the other hand, condition (C2) ensures the existence of a Lyapunov function that is strictly positive everywhere except at $p^{\dagger}$ and decreases along the flow of (23) ([Sas13]). Either condition guarantees the convergence of the trajectories of (23) to $p^{\dagger}$.
Let $D^{\dagger}$ denote the domain of attraction of $p^{\dagger}$ for the dynamical system (23). From converse Lyapunov theorem [Sch64], we know that there exists a continuously differentiable function $\bar{V}:D^{\dagger}\rightarrow\mathbb{R}_{+}$ such that $\bar{V}(p^{\dagger})=0$, $\bar{V}(p)>0$ for all $p\in D^{\dagger}\backslash\{p^{\dagger}\}$ and $\bar{V}(p)\rightarrow\infty$ as $p\rightarrow\textsf{boundary}(D^{\dagger})$. For any $r>0$, define $\bar{V}_{r}=\{p\in\textsf{dom}(\bar{V}):\bar{V}(p)\leqslant r\}$ to be a sub-level set of $\bar{V}$. There exists $0<\bar{r}^{\prime}<\bar{r}$ such that $\bar{V}_{\bar{r}^{\prime}}\subsetneq\mathcal{B}_{\bar{r}^{\prime}}(p^{\dagger})\subsetneq\mathcal{B}_{\bar{r}}(p^{\dagger})\subsetneq\bar{V}_{\bar{r}}$. Additionally, define $\ell_{0}=0,\ell_{k}=\sum_{i=1}^{k}\beta_{i}$ and $L_{k}=\ell_{n(k)}$ where $n(0)=0$, and
@@
To ensure that $\bar{\textbf{p}}^{(k)}(L_{k})\in\textsf{dom}(\bar{V})$ for $k>0$, we show that for an appropriate choice of $T$ in (24), $p_{L_{k}}\in\textsf{int}(D^{\dagger})$ for every $k\in\mathbb{N}$. From [BP18, Theorem IV.1], we know that there exists $K>0$ such that for all $k\in\mathbb{N},$
@@
Define $\hat{p}:\mathbb{R}_{+}\rightarrow\mathbb{R}$ such that, for every $k\in\mathbb{N}$, $\hat{p}(\ell_{k})=p_{k}$ with linear interpolation on $[\ell_{k},\ell_{k+1}]$. Using the standard approximation arguments from [Bor09, Chapter 6], it holds that333 For any $T\geqslant\bar{k}$ and $\delta>0$, there exists $k(\delta)$ such that $\hat{p}(\ell_{k(\delta)}+\cdot)$ form a “$(T,\delta)$” perturbation (cf. [Bor97]) of (23).
@@
Using (21) and Assumption 3.1, we conclude that RHS in the above equation goes to zero as $k\rightarrow\infty$. Finally, using (26), (3.2) and [Bor97, Lemma 2.1] we conclude that $p_{k}\rightarrow p^{\dagger}$ as $k\rightarrow\infty$. ∎

### v2

@@ 1 Introduction
Incentive mechanisms play a crucial role in many societal systems, where the outcomes are governed by the selfish interactions of a large number users (or algorithms on their behalf) and the system operator. The outcome of such strategic interactions, characterized by the Nash equilibrium, is often sub-optimal because of the fact that individual agents often do not account for the externality of their actions (i.e., how their actions affect the costs of others), when minimizing their own costs. An important way to address the sub-optimality is to provide agents with incentives that align their individual goal of cost minimization with the goal of minimizing the total cost of the societal system[Levin, 1985, Ho et al., 1982, Başar, 1984].
The system operator faces two key challenges when designing such incentive mechanisms. First, due to the system’s scale and privacy concerns, the designer often has limited information about individual agents’ costs and the cost function of the entire system. Second, the system operator often needs to consider the learning behavior of agents, who repeatedly update their strategies in response to the incentive mechanism, especially when the physical system experiences a random shock and players are trying to reach a new equilibrium [Barrera and Garcia, 2014, Como and Maggistro, 2021, Maheshwari et al., 2021]. As a result, the designer cannot infer the equilibrium to predict agents’ behavior, making incentive design a challenging task. This leads to the following question:
@@
We provide sufficient conditions on the underlying game such that the set of fixed points is a singleton, making the socially optimal incentive mechanism unique (Proposition 3.1). We also characterize sufficient conditions that guarantee the convergence of the coupled dynamical system (Proposition 3.3). Since the convergent strategy profile and incentive mechanism correspond to a fixed point that is also socially optimal, these sufficient conditions ensure that the dynamic incentive mechanism eventually induces a socially optimal outcome in the long run. To prove this result, we leverage the timescale separation between the strategy update and the incentive update, using results from the theory of two-timescale dynamical systems [Borkar, 1997]. This theory allows us to analyze the convergence of the coupled dynamical system in two steps by decoupling the convergence of the strategy update and the incentive mechanism update. First, the convergence of the strategy update (on the faster timescale) is analyzed by viewing the incentive mechanism (on the slower timescale) as static. In our work, we off-load this convergence to the rich literature on learning in games (e.g., [Fudenberg and Levine, 1998, Sandholm, 2010, Monderer and Shapley, 1996]). Second, the convergence of the incentive mechanism update is analyzed via the associated continuous-time dynamical system, evaluated at the converged value of the strategy update (i.e., Nash equilibrium). The sufficient conditions developed in Proposition 3.3 are built on results on the convergence of non-linear dynamical systems from cooperative dynamical systems theory [Hirsch, 1985] and Lyapunov-based methods [Sastry, 2013].
@@ 1.1 Related Works
Two-timescale Learning Dynamics: Learning dynamics where incentives are updated at a slower timescale than the strategy of agents have been studied in [Mojica-Nava et al., 2022, Chen et al., 2023, Poveda et al., 2017, Ochoa and Poveda, 2022, Ratliff and Fiez, 2020, Liu et al., 2021, Li et al., 2023b]. Specifically, [Mojica-Nava et al., 2022] studies Stackelberg games with a single leader and a population of followers. In their approach the leader deploys gradient-based update and followers update their strategies using replicator dynamics. [Chen et al., 2023] and [Poveda et al., 2017] consider incentive design in affine congestion games, where the incentives are updated using a distributed version of gradient descent. [Ochoa and Poveda, 2022] studies incentive design to control traffic congestion on a single highway using gradient based incentive updates. [Ratliff and Fiez, 2020] studies the problem of incentive design while learning the cost functions of agents. The authors assume that both the cost functions and incentive policies are linearly parameterized, and the incentive updates rely on the knowledge of agents’ strategy update rules instead of just the current strategy as in our setting. Additionally, [Liu et al., 2021, Li et al., 2023b] propose a two-timescale discrete-time learning dynamics, where agents adopt a mirror descent-based strategy update, and the social planner updates the incentive parameter according to a gradient based method. The convergence of these gradient-based learning dynamics relies on the assumption that the social cost given agents’ equilibrium strategy is convex in the incentive parameter. However, the convexity assumption is restrictive as highlighted by an example in Appendix A.
@@
Same-timescale Learning Dynamics: The problem of steering non-cooperative agents to a desired Nash equilibrium by considering an incentive update which runs on the same timescale as strategy updates is studied in [Shakarami et al., 2022a, Shakarami et al., 2022b, Zhang et al., 2023]. Particularly, in [Shakarami et al., 2022a] the authors study such updates in a specific setting of quadratic aggregative games. In [Shakarami et al., 2022b], the authors consider a setup where the agents’ cost is only dependent on their own action and the price signal which is provided by operator which is used by agents based on their individual reliability over the operator. The aim is to design this price signal so that the reliability on the price signal is retained while steering the equilibrium to the desired point. In [Zhang et al., 2023], the authors consider the problem of steering no-regret learning agents to optimal equilibrium but the computation of incentive mechanism requires solving an optimization problem at each time step.
@@
Learning in Stackelberg Games: Our work is also related to the literature on learning in Stackelberg games, where the planner often has limited information about the game between agents and need to design optimal mechanism by dynamically receiving feedback from agents’ responses (for instance, refer to [Blum et al., 2014, Letchford et al., 2009, Peng et al., 2019, Bai et al., 2021, Cui et al., 2024, Jalota et al., 2022, Maheshwari et al., 2023]). Typically, this line of research posit structural assumptions on the game between followers (e.g., finite action space of agents, or linearly parametrized utility function of agents)[Blum et al., 2014, Letchford et al., 2009, Peng et al., 2019, Bai et al., 2021, Cui et al., 2024, Jalota et al., 2022], or only ensures convergence to a locally optimal solution [Maheshwari et al., 2023].
@@ 2.1.1 Atomic Games
Under Assumption 2.1, it is known that, for any $p\in\mathbb{R}^{|\mathcal{I}|}$, the Nash equilibrium $x^{\ast}(p)$ exists, is unique and is Lipscthiz continuous in $p$ [Dafermos, 1988]. Moreover, the socially optimal strategy $x^{\dagger}$ is also unique.
@@ 2.1.2 Non-atomic Games
Under Assumption 2.2, Nash equilibrium $x^{\ast}(p)$ exists, is unique and Lipscthiz continuous for every $p\in\mathbb{R}^{|\tilde{\mathcal{I}}|}$ [Sandholm, 2010, Dafermos, 1988]. Moreover, the socially optimal strategy distribution is also unique.
@@ 2.2 Coupled Strategy and Incentive Update
    Equilibrium update ([Cui et al., 2024, Jalota et al., 2022]): The strategy update incorporates a Nash equilibrium strategy profile with respect to the incentive mechanism in step $k$:
@@
    Best response update ([Fudenberg and Levine, 1998, Hofbauer and Sandholm, 2002]): The strategy update incorporates a best response strategy with respect to the strategy and the incentive mechanism in step $k$:
@@
    Gradient-based update ([Littlestone and Warmuth, 1994, Laraki and Mertikopoulos, 2013, Sandholm, 2010]): The strategy update incorporates a smoothed strategy update with respect to the strategy and the incentive mechanism in step $k$:
@@
    where $z_{i}(x_{k},p_{k})=x_{k}-\eta\nabla_{x_{i}}c_{i}(x_{k},p_{k})$, $\eta$ is step size, and $h(\cdot),\tilde{h}(\cdot)$ are regularizers. If $h(\cdot)$ is a quadratic function then the update becomes projected gradient descent [Mertikopoulos and Zhou, 2019]. Furthermore, if $\tilde{h}(\cdot)$ is the entropy function then the update becomes a perturbed best-response update[Sandholm, 2010].
@@
Our externality-based incentive update contrasts with the gradient-based incentive update considered in the literature [Liu et al., 2021]. Such a gradient-based incentive update typically requires that the equilibrium social cost function $\Phi(x^{\ast}(p))$ (resp. $\tilde{\Phi}(\tilde{x}^{\ast}(\tilde{p}))$) be strongly convex in the incentive mechanism $p$ (resp. $\tilde{p}$). However, this requirement is too restrictive and is not satisfied even in simple games due to the fact that the equilibrium strategy vector is typically non-convex in the incentive mechanism. In Appendix A, we provide an example of a two-link routing game, where the equilibrium social cost function is non-convex in the incentive mechanism, and the gradient-based learning dynamics do not converge.
@@ 3.1 Fixed point analysis
First, we show that $P^{\dagger}$ is non-empty. That is, there exists $p^{\dagger}$ such that $e(x^{\ast}(p^{\dagger}))=p^{\dagger}$. Define a function $\theta(p)=e(x^{\ast}(p))$. We note that Assumption 2.1 guarantees that $\theta$ is well defined. Thus the problem is reduced to showing existence of fixed point of $\theta(\cdot)$, for which we use the Schauder fixed point theorem [Smart, 1980].
@@
where $J(x,p)$ is the game Jacobian defined as $J_{i}(x,p)=\nabla_{{x}_{i}}\ell_{i}(x)+p_{i}$ for every $i\in\mathcal{I}$. Next, under Assumption 2.1, the variational inequality characterization of Nash equilibrium [Facchinei and Pang, 2007], we know that $x^{\ast}(p^{\dagger})$ is a Nash equilibrium if and only if
@@
Under Assumption 2.2, the variational inequality characterization of Nash equilibrium for non-atomic games [Sandholm, 2010], we know that ${\tilde{x}^{\ast}}(\tilde{p}^{\dagger})$ is a Nash equilibrium if and only if
@@ 3.2 Convergence to optimal incentive mechanism
In Assumption 3.1, (i) is a standard assumption on step-sizes that allows us to analyze the convergence of the discrete-time learning updates via a continuous-time differential equation [Borkar, 1997]. Additionally, (ii) assumes that the incentive update occurs at a slower timescale compared to the update of strategies.
In view of Assumption 3.1-(ii), it is necessary that the strategy updates converge if the incentives are held fixed. This condition holds true for a variety of updates commonly studied in learning in games literature. For instance, the equilibrium update (cf. (4)) converges to Nash equilibrium in one-step. Similarly, the best response update (cf. (5)) converges in zero sum games [Harris, 1998], potential games [Swenson et al., 2018], and dominance solvable games [Nachbar, 1990]. The gradient-based update (cf. (6)) converge in concave games [Rosen, 1965]. Therefore, in this work we do not focus on studying the convergence of strategy updates with fixed incentive mechanisms and we pick any off-the-shelf convergent strategy update. Our goal is to characterize conditions under which the coupled strategy and incentive updates ($x$-update)-($p$-update) and ($\tilde{x}$-update)-($\tilde{p}$-update) converge. Therefore, we impose the following assumption which is sufficient to ensure convergence of strategy updates ($x$-update) and ($\tilde{x}$-update) with fixed incentive mechanisms. Particularly, we assume that if the incentive mechanism is fixed, then the corresponding continuous-time dynamical system converges to the Nash equilibrium associated with that incentive mechanism.
@@
Using the standard approximation arguments from [Borkar, 2009] it is evident that Assumption 3.1 and 3.2 are sufficient to ensure convergence of strategy updates ($x$-update) and ($\tilde{x}$-update) with fixed incentive mechanisms.
@@
satisfy [^2] The results in this paper hold even if the updates ($x$-update)-($p$-update) and ($\tilde{x}$-update)-($\tilde{p}$-update) are perturbed with martingale difference noise [Borkar and Pattathil, 2018].
@@
Under Assumption 3.1, the two-timescale approximation theory [Borkar, 1997, Borkar and Pattathil, 2018] suggests that the strategy update ($x$-update) is a fast transient while the incentive update ($p$-update) is a slow component. This separation in timescales allow us to study the convergence of these updates in two stages. First, we study the convergence of fast strategy updates, for every fixed value of incentive. Second, we study the convergence of slow incentive updates, assuming that the fast strategy updates have converged to the equilibrium.
@@
Since $\sup_{k\in\mathbb{N}}(\|x_{k}\|+\|p_{k}\|)<+\infty$ and $\lim_{k\rightarrow\infty}\beta_{k}/\gamma_{k}=0$ (cf. Assumption 3.1), the term $\frac{\beta_{k}}{\gamma_{k}}\left(e(x_{k})-p_{k}\right)$ in (20) goes to zero as $k\rightarrow\infty$. Consequently, leveraging the standard approximation arguments [Borkar, 2009, Lemma 1, Section 2.2], we conclude that the asymptotic behavior of the updates in (20) is same as that of the following dynamical system
@@
Note that $p^{\dagger}$ is the fixed point of the trajectories of dynamical system (23) (cf. Proposition 3.1). Conditions (C1) and (C2) in Proposition 3.3 are based on results from non-linear dynamical systems which ensure convergence of (23). In particular, (C1)-(i) (resp. (C1)-(ii)) builds on the cooperative dynamical systems theory [Hirsch, 1985], which ensures that $\mathbb{R}_{+}^{|\mathcal{I}|}$ (resp. $\mathbb{R}_{-}^{|\mathcal{I}|}$) is positively invariant for (23) and $p^{\dagger}\in\mathbb{R}_{+}^{|\mathcal{I}|}$ (resp. $p^{\dagger}\in\mathbb{R}_{-}^{|\mathcal{I}|}$) is asymptotically stable. On the other hand, condition (C2) ensures the existence of a Lyapunov function that is strictly positive everywhere except at $p^{\dagger}$ and decreases along the flow of (23) ([Sastry, 2013]). Either condition guarantees the convergence of the trajectories of (23) to $p^{\dagger}$.
Let $D^{\dagger}$ denote the domain of attraction of $p^{\dagger}$ for the dynamical system (23). From converse Lyapunov theorem [Scheidegger, 1964], we know that there exists a continuously differentiable function $\bar{V}:D^{\dagger}\rightarrow\mathbb{R}_{+}$ such that $\bar{V}(p^{\dagger})=0$, $\bar{V}(p)>0$ for all $p\in D^{\dagger}\backslash\{p^{\dagger}\}$ and $\bar{V}(p)\rightarrow\infty$ as $p\rightarrow\textsf{boundary}(D^{\dagger})$. For any $r>0$, define $\bar{V}_{r}=\{p\in\textsf{dom}(\bar{V}):\bar{V}(p)\leqslant r\}$ to be a sub-level set of $\bar{V}$. There exists $0<\bar{r}^{\prime}<\bar{r}$ such that $\bar{V}_{\bar{r}^{\prime}}\subsetneq\mathcal{B}_{\bar{r}^{\prime}}(p^{\dagger})\subsetneq\mathcal{B}_{\bar{r}}(p^{\dagger})\subsetneq\bar{V}_{\bar{r}}$. Additionally, define $\ell_{0}=0,\ell_{k}=\sum_{i=1}^{k}\beta_{i}$ and $L_{k}=\ell_{n(k)}$ where $n(0)=0$, and
@@
To ensure that $\bar{\textbf{p}}^{(k)}(L_{k})\in\textsf{dom}(\bar{V})$ for $k>0$, we show that for an appropriate choice of $T$ in (24), $p_{L_{k}}\in\textsf{int}(D^{\dagger})$ for every $k\in\mathbb{N}$. From [Borkar and Pattathil, 2018, Theorem IV.1], we know that there exists $K>0$ such that for all $k\in\mathbb{N},$
@@
Define $\hat{p}:\mathbb{R}_{+}\rightarrow\mathbb{R}$ such that, for every $k\in\mathbb{N}$, $\hat{p}(\ell_{k})=p_{k}$ with linear interpolation on $[\ell_{k},\ell_{k+1}]$. Using the standard approximation arguments from [Borkar, 2009, Chapter 6], it holds that333 For any $T\geqslant\bar{k}$ and $\delta>0$, there exists $k(\delta)$ such that $\hat{p}(\ell_{k(\delta)}+\cdot)$ form a “$(T,\delta)$” perturbation (cf. [Borkar, 1997]) of (23).
@@
Using (21) and Assumption 3.1, we conclude that RHS in the above equation goes to zero as $k\rightarrow\infty$. Finally, using (26), (3.2) and [Borkar, 1997, Lemma 2.1] we conclude that $p_{k}\rightarrow p^{\dagger}$ as $k\rightarrow\infty$. ∎